# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v2

@@ II.1 Modal formulation of the scattering problem
In general, the electromagnetic resonances of a 3D, lossy and dispersive resonator are most rigorously treated as QNMs Sauvan et al. (2022). We consider a system comprising a lossy resonator, contained in some compact region $\Omega_{\mathrm{res}}\subset\mathbb{R}^{3}$ and embedded in some arbitrary linear material background described by a relative permittivity $\varepsilon_{\mathrm{b}}(\bm{r},\omega)$ and relative permeability $\mu_{\mathrm{b}}(\bm{r},\omega)$, while we denote the relative material parameters of the whole system by $\varepsilon(\bm{r},\omega)$ and $\mu(\bm{r},\omega)$. In general, the background does not need to be homogeneous and could, for example, contain a substrate or waveguides, as schematized in Fig. 1(a). The QNM electric and magnetic fields, ${\bm{E}}_{n}$ and ${\bm{H}}_{n}$ respectively, are defined as the time-harmonic solutions of the source-free Maxwell’s equations,
@@
subject to an appropriate outgoing-wave boundary condition (note that we have assumed a time dependence of the form $\mathrm{e}^{-\mathrm{i}\omega t}$). Alternatively, the QNM problem in Eq. (II.1) can be formulated as an equation for the electric field only:
×$\nabla\times\left(\mu^{-1}(\bm{r},\tilde{\omega}_{n})\nabla\times\bm{E}_{n}\right)-\frac{\tilde{\omega}_{n}^{2}}{c^{2}}\varepsilon(\bm{r},\tilde{\omega}_{n})\bm{E}_{n}=0\ .$ (2)
In the case of a homogeneous background, one can use the Silver-Müller radiation condition, but in a more complex background the condition needs to be adjusted (see, for example, Refs. Kristensen, de Lasson, and Gregersen (2014); Kristensen et al. (2017)).
The QNM eigenfrequencies $\tilde{\omega}_{n}=\omega_{n}-\mathrm{i}\gamma_{n}$ are generally complex in character by virtue of losses  [^1] Notably however, static/longitudinal modes with $\tilde{\omega}_{\mu}=0\in\mathbb{R}$ can occur, and the significance of their non-resonant contribution in QNM theories has been investigated in the literature Sauvan (2021); Besbes and Sauvan (2022).. Due to their exponential decay in time ($\gamma_{n}<0$), the QNMs exponentially diverge in space at $|\bm{r}|\rightarrow\infty$ Sauvan et al. (2022). Moreover, since they satisfy an outgoing-wave condition, the QNMs can only be used to expand scattered fields. It is therefore natural to adopt a background-field formulation which splits the total field into $\bm{E}(\bm{r},\omega)=\bm{E}_{\mathrm{b}}(\bm{r},\omega)+\bm{E}_{\mathrm{sca}}(\bm{r},\omega)$. Here, the background or incident field $\bm{E}_{\mathrm{b}}(\bm{r},\omega)$ satisfies Maxwell’s equations in the background medium,
×$\nabla\times\left(\mu^{-1}_{\mathrm{b}}\nabla\times\bm{E}_{\mathrm{b}}\right)-\frac{\omega^{2}}{c^{2}}\varepsilon_{\mathrm{b}}\bm{E}_{\mathrm{b}}=\mathrm{i}\omega\mu_{0}\bm{j}_{\mathrm{b}}\ ,$ (3)
@@
[…]
@@
where $\Delta\varepsilon=\varepsilon-\varepsilon_{\mathrm{b}}$, $\Delta\mu^{-1}=\mu^{-1}-\mu^{-1}_{\mathrm{b}}$ [$\mathrm{supp}(\Delta\varepsilon(\cdot,\omega)/\Delta\mu^{-1}(\cdot,\omega))\subseteq\Omega_{\mathrm{res}}$], and the same outgoing-wave condition as for the QNMs.
With regard to the practical calculation of QNMs and the formulation of a modal theory, two problems arise that need to be addressed. Considering their calculation, an outgoing-wave condition is usually imposed as an asymptotic one for the fields at $|\bm{r}|\rightarrow\infty$, while in numerical calculations the domain must be truncated. Using the QNMs in an expansion of the scattered field $\bm{E}_{\mathrm{sca}}$ at real frequencies $\omega$, it is apparent that this expansion cannot be valid at positions far away from the resonator due to their spatial divergence. In fact, it is known that the QNMs can only be used to expand the scattered field and the system’s Green’s function for positions inside the resonator, where they appear to form a complete set, provided that the resonator is defined by a discontinuity in the refractive index in space Leung, Liu, and Young (1994); Leung and Pang (1996); Lee, Leung, and Pang (1999). However, even this is only true for trivial homogeneous backgrounds, and becomes generally invalid for more complex background geometries, such as the one sketched in Fig. 1, due to branch cuts of the Green’s function Yan, Faggiani, and Lalanne (2018); Sauvan et al. (2022). A general QNM expansion of the scattered field $\bm{E}_{\mathrm{sca}}(\bm{r},\omega)$ would thus have the form
@@
Figure 1: Concept of the background-field formulation and the PML-based truncation of the domain used for QNM calculation, regularization and quantization. (a) Schematic of an example background, described by material parameters $(\varepsilon_{\mathrm{b}},\mu_{\mathrm{b}})$, in which Maxwell’s equations have to be satisfied by the background/incident field $\bm{E}_{\mathrm{b}}$ in the infinite space. The background can be non-trivial and contain substrates or waveguides. (b) The scattered-field problem in which Maxwell’s equations have to be satisfied by $\bm{E}_{\mathrm{sca}}$ in the full geometry, including the resonator with compact domain $\Omega_{\mathrm{res}}$ and containing the effective current density arising from $\bm{E}_{\mathrm{b}}$. Suitable outgoing-wave conditions in the infinite space are applied. (c) The corresponding problem for $\bm{E}_{\mathrm{sca}}$ with PMLs imposed, in which the infinite space is truncated to a compact domain $\Omega$. Using a complex coordinate transformation, which leaves the inner domain $\Omega\setminus\Omega_{\mathrm{PML}}$ invariant and causes outgoing waves to be damped in $\Omega_{\mathrm{PML}}$, is equivalent to the introduction of an effective material $(\tilde{\varepsilon},\tilde{\mu})$. The damping of the outgoing waves effectively converts the outgoing-wave condition to a Dirichlet condition (perfect electrical conductor) at infinity, and due to the exponential nature of the damping, this boundary condition can very well be approximated using a finite truncated domain $\Omega$ and imposing the condition at $\partial\Omega$.
PMLs were originally introduced in computational electrodynamics in order to simulate open boundaries Berenger (1994). They are defined as material layers that surround a finite domain such that the dynamics therein are unaffected, i.e., outgoing waves are transmitted, without reflection, into the PML region $\Omega_{\mathrm{PML}}$ and are exponentially damped within them. The outgoing wave condition is thus effectively transformed into a Dirichlet (i.e., perfect electrical conductor) boundary condition at $|\bm{r}|\rightarrow\infty$. Due to the fast exponential damping of the waves, it is a good approximation to consider PMLs with finite thickness only, so that the infinite space $\mathbb{R}^{3}$ has been effectively mapped to a finite domain $\Omega$, in which we solve the scattered-field problem. This situation is sketched in Fig. 1(c). We denote the transformed material parameters of the mapped system by $(\tilde{\varepsilon}(\bm{r},\omega),\tilde{\mu}(\bm{r},\omega))$, whereas the inner domain $\Omega\setminus\Omega_{\mathrm{PML}}$ is not affected by this transformation. When using PMLs, however, there are some limitations. In particular, the material parameters are required to be analytic functions in the directions perpendicular to the PML boundaries (connected to the idea of PMLs as complex coordinate transformations), implying that the PMLs must be chosen in such a manner that waveguides enter with an angle of $90^{\circ}$, and periodic photonic crystal environments cannot be treated without further considerations Oskooi et al. (2008).
@@
By using the correspondence between coordinate transformations and effective materials in electrodynamics Ward and Pendry (1996), Maxwell’s equations with the material parameters $(\varepsilon,\mu)$ in the coordinates $\tilde{x}_{i}$ can be rewritten in the real coordinates $x_{i}$ with effective materials $(\tilde{\varepsilon},\tilde{\mu})$ using the transformations
@@
The solutions of this transformed and PML-truncated eigenvalue problem are a discrete set of modes $\tilde{\bm{E}}_{n}(\bm{r})$ that can be grouped into QNM-like modes $\tilde{\bm{E}}_{n}^{\mathrm{QNM}}(\bm{r})$, which are good approximations of the true QNMs of the system and do not depend on the parameters of the PMLs, and so-called numerical or PML modes $\tilde{\bm{E}}_{n}^{\mathrm{PML}}(\bm{r})$ that depend sensitively on the PML parameters and which originate from the continuous non-resonant contribution in Eq. (5). Hence, by introducing finite-thickness PMLs, this background is effectively mapped to a discrete set of modes Olyslager (2004). Since the operator in Eq. (II.1) is not hermitian, the completeness of these eigenmodes is not always guaranteed and we must further assume the absence of exceptional points where two eigenvectors coalesce. In this case, the PML-regularized scattered field $\tilde{\bm{E}}_{\mathrm{sca}}$ can be expanded in this discrete set of modes,
@@
where, henceforth, we shall not distinguish between the QNM-like and PML modes. Mathematically, they arise as a common set of solutions to the eigenvalue problem in Eq. (II.1), and should be treated on an equal footing for the purpose of further analysis.
In the context of non-dispersive magnetic permeabilities $\mu=\mu(\bm{r})$, Yan et al. Yan, Faggiani, and Lalanne (2018) have developed an approach for rigorous QNM analysis of lossy resonators, relying on an auxiliary-field eigenvalue technique with PML regularization, within the framework of a finite-element analysis that enables efficient numerical calculations. Whilst the quantization scheme that we will discuss in § III can in principle be used for arbitrary dispersions of $(\tilde{\varepsilon},\tilde{\mu})$, the approach of Yan et al. facilitates practical calculations in the important special cases of dielectric and plasmonic cQED, and formulae obtained within it can be used in the quantized theory for these cases as well (such as, for instance, the analytic excitation coefficients of the modes for an external driving field $\bm{E}_{\mathrm{b}}$). As such, we will briefly discuss the key results of this approach in the next subsection.
II.2 Auxiliary-field eigenvalue technique
The non-linear nature of the eigenvalue problem in Eq. (II.1) for dispersive materials renders its numerical solution difficult. The introduction of auxiliary fields in accordance with Yan et al. Yan, Faggiani, and Lalanne (2018) aims to linearize the eigenvalue problem so that more conventional and efficient solvers can be leveraged. The permittivity distribution $\varepsilon(\bm{r},\omega)$ is described using a position-dependent $N$-pole Drude-Lorentz model,
$\varepsilon(\bm{r},\omega)=\varepsilon_{\infty}(\bm{r})-\varepsilon_{\infty}(\bm{r})\sum_{l=1}^{N}\frac{\omega_{\text{p},l}^{2}(\bm{r})}{\omega^{2}-\omega_{0,l}^{2}(\bm{r})+\mathrm{i}\gamma_{l}(\bm{r})\omega},$
in which $\gamma_{l}(\bm{r})$, $\omega_{\text{p},l}(\bm{r})$ and $\omega_{0,l}(\bm{r})$ are damping rates, plasma and resonance frequencies pertaining to the $l$th pole, respectively. Dispersive PMLs like the one defined in Eq. (6) destroy the form of the Drude-Lorentz permittivity in $\Omega_{\mathrm{PML}}$ and it is therefore common to use non-dispersive PMLs, in which the $\omega$-dependence is replaced by a constant frequency $\omega_{\mathrm{PML}}$ located within the frequency range of interest. These non-dispersive PMLs have the disadvantage that modes with lower frequency are less rapidly attenuated, thereby compromising the accuracy of the simulation. However, alternative auxiliary field methods can also be derived for arbitrary, rational dispersive permittivities (including dispersive PMLs), and studies with dispersive PMLs have been performed Lalanne et al. (2019); Demésy et al. (2023). For the non-dispersive PMLs, only $\varepsilon_{\infty}(\bm{r})$ in Eq. (II.2) needs to be transformed to $\tilde{\varepsilon}_{\infty}(\bm{r})$, and auxiliary polarizations $\tilde{\bm{P}}_{l}$ and current densities $\tilde{\bm{J}}_{l}$ are defined for each pole, via
$\tilde{\bm{P}}_{l}=$ $-\frac{\varepsilon_{0}\tilde{\varepsilon}_{\infty}(\bm{r})\omega_{\text{p},l}^{2}(\bm{r})}{\omega^{2}-\omega_{0,l}^{2}(\bm{r})+\mathrm{i}\omega\gamma_{l}(\bm{r})}\tilde{\bm{E}},\quad\tilde{\bm{J}}_{l}=-\mathrm{i}\omega\tilde{\bm{P}}_{l}\,.$ (13)
Treating the simple case of a single Drude-Lorentz pole $(N=1)$, we use these auxiliary fields to reformulate Eq. (II.1) as a linear eigenvalue problem, ^$\hat{\mathcal{H}}\tilde{\bm{\psi}}_{n}={\tilde{\omega}}_{n}\tilde{\bm{\psi}}_{n}$, for an augmented eigenvector $\tilde{\bm{\psi}}_{n}=[\tilde{\bm{H}}_{n},\tilde{\bm{E}}_{n},\tilde{\bm{P}}_{n},\tilde{\bm{J}}_{n}]^{\text{T}}$, in which
^$\hat{\mathcal{H}}=\begin{bmatrix}0&-\mathrm{i}(\mu_{0}\tilde{\mu})^{-1}\nabla\times&0&0\\ \mathrm{i}(\varepsilon_{0}\tilde{\varepsilon}_{\infty}^{-1})\nabla\times&0&0&-\mathrm{i}(\varepsilon_{0}\tilde{\varepsilon}_{\infty})^{-1}\\ 0&0&0&\mathrm{i}\\ 0&\mathrm{i}\varepsilon_{0}\tilde{\varepsilon}_{\infty}\omega_{\text{p}}^{2}&-\mathrm{i}\omega_{0}^{2}&-\mathrm{i}\gamma\end{bmatrix}\,.$ (14)
The generalization to an arbitrary number $N$ of poles is straightforward.
From the unconjugated form of the Lorentz reciprocity theorem, applied to the augmented system in Eq. (14), an orthogonality relation for the modes can be derived:
$\int_{\Omega}\mathrm{d}^{3}r\,\bigg{(}\varepsilon_{0}\tilde{\varepsilon}_{\infty}(\bm{r})\tilde{\bm{\bm{E}}}_{n}\cdot\tilde{\bm{E}}_{n^{\prime}}-\mu_{0}\tilde{\mu}(\bm{r})\tilde{\bm{H}}_{n}\cdot\tilde{\bm{H}}_{n^{\prime}}$ (15)
$+\frac{\omega_{0}^{2}(\bm{r})}{\varepsilon_{0}\tilde{\varepsilon}_{\infty}\omega_{\text{p}}^{2}(\bm{r})}\tilde{\bm{P}}_{n}\cdot\tilde{\bm{P}}_{n^{\prime}}-\frac{1}{\varepsilon_{0}\tilde{\varepsilon}_{\infty}(\bm{r})\omega_{\text{p}}^{2}(\bm{r})}\tilde{\bm{J}}_{n}\cdot\tilde{\bm{J}}_{n^{\prime}}\bigg{)}=\delta_{nn^{\prime}}\,,$
and their normalization condition can be readily written as
@@ II.2 Auxiliary-field eigenvalue technique
$=1\,.$ (16)
It should be emphasized that the volume integration in these relations is performed over the entire computational domain $\Omega$, including the PML regions $\Omega_{\mathrm{PML}}$, yet the normalization of the QNMs does not depend on the PML parameters Sauvan et al. (2022). Although the orthogonality relation in this form can only be derived within the auxiliary-field eigenvalue approach, the normalization can also be used for materials with arbitrary dispersion and permeability, for which we would have to replace $\tilde{\mu}\rightarrow\partial_{\omega}[\omega\tilde{\mu}(\bm{r},\omega)]$.
By defining the matrix
$\hat{\mathcal{D}}=\mathrm{diag}\left[-\mu_{0}\tilde{\mu},\varepsilon_{0}\tilde{\varepsilon}_{\infty},\frac{\omega_{0}^{2}}{\varepsilon_{0}\tilde{\varepsilon}_{\infty}\omega_{\mathrm{p}}^{2}},\frac{-1}{\varepsilon_{0}\tilde{\varepsilon}_{\infty}\omega_{\mathrm{p}}^{2}}\right]\,,$ (17)
corresponding to the orthogonality relation in Eq. (15), the completeness relation can be written as
$\sum_{n}\tilde{\bm{\psi}}_{n}(\bm{r}^{\prime})\otimes\tilde{\bm{\psi}}_{n}^{\mathrm{T}}(\bm{r})\hat{\mathcal{D}}(\bm{r})=\hat{\bm{I}}\delta(\bm{r}-\bm{r}^{\prime})\,,$ (18)
in which $\tilde{\bm{\psi}}_{n}^{\mathrm{T}}(\bm{r})\hat{\mathcal{D}}$ are the left eigenvectors of $\hat{\mathcal{H}}$ Yan, Faggiani, and Lalanne (2018). Invoking the orthogonality and completeness properties, as well as the scattered-field formulation, we can derive analytic expressions for the modal excitation coefficients in Eq. (II.1):
@@
$-\frac{\tilde{\omega}_{n}}{\tilde{\omega}_{n}-\omega}\mu_{0}$ $\langle\tilde{\bm{H}}^{*}_{n}|\Delta\mu(\bm{r})|\bm{H}_{\mathrm{b}}\rangle_{L^{2}(\Omega_{\mathrm{res}})}\,,$ (19)
in which $\langle\cdot|\cdot\rangle_{L^{2}(\Omega_{\mathrm{res}})}$ is the standard complex $L^{2}$-inner product over $\Omega_{\mathrm{res}}$. With these formulae, we can calculate the excitation coefficients in the frequency domain for arbitrary incident fields $\bm{E}_{\mathrm{b}}(\bm{r},\omega)$ through simple overlap integrals. A modal expansion as in Eq. (II.1) is also possible in the time domain with coefficients $\xi_{n}(t)$ related to $\alpha_{n}(\omega)$ via a Fourier transform and satisfying the equation
@@
$-\varepsilon_{0}\langle\tilde{\bm{E}}_{n}^{*}|\varepsilon_{\mathrm{b}}-\varepsilon_{\infty}|\frac{\mathrm{d}}{\mathrm{d}t}\bm{E}_{\mathrm{b}}(t)\rangle-\mathrm{i}\tilde{\omega}_{n}\mu_{0}\langle\tilde{\bm{H}}_{n}^{*}|\Delta\mu|\bm{H}_{\mathrm{b}}(t)\rangle$
@@
It is worth noting that although this auxiliary-field eigenvalue approach was originally introduced for systems with non-dispersive magnetic permeabilities $\mu(\bm{r})$, its generalization for media with strongly frequency-dependent permeabilities (as relevant for magnonic cQED) may be feasible via the introduction of suitable auxiliary fields for the magnetic response. In the following, we assume that such a complete set of suitably regularized and normalized modes can be ascertained, and explore their quantization for arbitrary, linear, 3D, magnetodielectric systems.

### v3

@@ II Classical QNMs: Regularization, Normalization and Completeness Based on Stretched Coordinates
We consider a system comprising a lossy resonator, contained in some compact region $\Omega_{\mathrm{res}}\subset\mathbb{R}^{3}$ and embedded in some arbitrary linear material background described by a relative permittivity $\varepsilon_{\mathrm{b}}(\bm{r},\omega)$ and relative permeability $\mu_{\mathrm{b}}(\bm{r},\omega)$, while we denote the relative material parameters of the whole system by $\varepsilon(\bm{r},\omega)$ and $\mu(\bm{r},\omega)$. In general, the background does not need to be homogeneous and could, for example, contain a substrate or waveguides, as schematized in Fig. 1(a). The QNM electric and magnetic fields, ${\bm{E}}_{n}$ and ${\bm{H}}_{n}$ respectively, are defined as the time-harmonic solutions of the source-free Maxwell’s equations,
@@
subject to an appropriate outgoing-wave boundary condition (note that we have assumed a time dependence of the form $\mathrm{e}^{-\mathrm{i}\omega t}$). Alternatively, the QNM problem in Eq. (II) can be formulated as an equation for the electric field only:
$\nabla\times\left(\mu^{-1}(\bm{r},\tilde{\omega}_{n})\nabla\times\bm{E}_{n}\right)-\frac{\tilde{\omega}_{n}^{2}}{c^{2}}\varepsilon(\bm{r},\tilde{\omega}_{n})\bm{E}_{n}=0\ .$ (2)
In the case of a homogeneous background, one can use the Silver-Müller radiation condition, but in a more complex background the condition needs to be adjusted (see, for example, Refs. kristensen2014calculation; kristensen2017theory).
The QNM eigenfrequencies $\tilde{\omega}_{n}=\omega_{n}-\mathrm{i}\gamma_{n}$ are generally complex in character by virtue of losses  [^1] Notably however, static/longitudinal modes with $\tilde{\omega}_{\mu}=0\in\mathbb{R}$ can occur, and the significance of their non-resonant contribution in QNM theories has been investigated in the literature sauvan2021quasinormal; besbes2022role.. Due to their exponential decay in time ($\gamma_{n}<0$), the QNMs exponentially diverge in space at $|\bm{r}|\rightarrow\infty$ sauvan2022norm. Moreover, since they satisfy an outgoing-wave condition, the QNMs can only be used to expand scattered fields. It is therefore natural to adopt a background-field formulation which splits the total field into $\bm{E}(\bm{r},\omega)=\bm{E}_{\mathrm{b}}(\bm{r},\omega)+\bm{E}_{\mathrm{sca}}(\bm{r},\omega)$. Here, the background or incident field $\bm{E}_{\mathrm{b}}(\bm{r},\omega)$ satisfies Maxwell’s equations in the background medium,
$\nabla\times\left(\mu^{-1}_{\mathrm{b}}\nabla\times\bm{E}_{\mathrm{b}}\right)-\frac{\omega^{2}}{c^{2}}\varepsilon_{\mathrm{b}}\bm{E}_{\mathrm{b}}=\mathrm{i}\omega\mu_{0}\bm{j}_{\mathrm{b}}\ ,$ (3)
@@
$\nabla\times$ $\left(\mu^{-1}\nabla\times\bm{E}_{\mathrm{sca}}\right)-\frac{\omega^{2}}{c^{2}}\varepsilon\bm{E}_{\mathrm{sca}}$
@@
where $\Delta\varepsilon=\varepsilon-\varepsilon_{\mathrm{b}}$, $\Delta\mu^{-1}=\mu^{-1}-\mu^{-1}_{\mathrm{b}}$ $\mathrm{supp}(\Delta\varepsilon(\cdot,\omega)/\Delta\mu^{-1}(\cdot,\omega))\subseteq\Omega_{\mathrm{res}}$], and the same outgoing-wave condition as for the QNMs.
With regard to the practical calculation of QNMs and the formulation of a modal theory, two problems arise that need to be addressed. Considering their calculation, an outgoing-wave condition is usually imposed as an asymptotic one for the fields at $|\bm{r}|\rightarrow\infty$, while in numerical calculations the domain must be truncated. Using the QNMs in an expansion of the scattered field $\bm{E}_{\mathrm{sca}}$ at real frequencies $\omega$, it is apparent that this expansion cannot be valid at positions far away from the resonator due to their spatial divergence. In fact, it is known that the QNMs can only be used to expand the scattered field and the system Green’s function for positions inside the resonator, where they appear to form a complete set, provided that the resonator is defined by a discontinuity in the refractive index in space leung1994completeness; leung1996dielectric; lee1999dyadic. However, even this is only true for trivial homogeneous backgrounds, and becomes generally invalid for more complex background geometries, such as the one sketched in Fig. 1, due to branch cuts of the Green’s function yan2018rigorous; sauvan2022norm (meaning that in such cases, the QNMs themselves generally do not form a complete set inside the resonator). A general QNM expansion of the scattered field $\bm{E}_{\mathrm{sca}}(\bm{r},\omega)$ would thus have the form
@@
[…]
PMLs were originally introduced in computational electrodynamics in order to simulate open boundaries berenger1994perfectly. They are defined as material layers that surround a finite domain such that the dynamics therein are unaffected, i.e., outgoing waves are transmitted, without reflection, into the PML region $\Omega_{\mathrm{PML}}$ and are exponentially damped within them. The outgoing wave condition is thus effectively transformed into a Dirichlet (i.e., perfect electrical conductor) boundary condition at $|\bm{r}|\rightarrow\infty$. Due to the fast exponential damping of the waves, it is a good approximation to consider PMLs with finite thickness only, so that the infinite space $\mathbb{R}^{3}$ has been effectively mapped to a finite domain $\Omega$, in which we solve the scattered-field problem. This situation is sketched in Fig. 1(c). We denote the transformed material parameters of the mapped system by $(\tilde{\varepsilon}(\bm{r},\omega),\tilde{\mu}(\bm{r},\omega))$, whereas the inner domain $\Omega\setminus\Omega_{\mathrm{PML}}$ is not affected by this transformation. When using PMLs, however, there are some limitations. In particular, the material parameters are required to be analytic functions in the directions perpendicular to the PML boundaries (connected to the idea of PMLs as complex coordinate transformations), implying that the PMLs must be chosen in such a manner that waveguides enter with an angle of $90^{\circ}$, and periodic photonic crystal environments cannot be treated without further considerations oskooi2008failure. In this case, it is believed that the QNMs are not complete inside the resonator sauvan2022norm, but notably, efforts to adapt QNM methodology for coupled cavity-waveguide systems have been made via the introduction of a non-local boundary condition kristensen2014calculation.
@@
[…]
@@
The solutions of this transformed and PML-truncated eigenvalue problem are a discrete set of modes $\tilde{\bm{E}}_{n}(\bm{r})$ that can be grouped into QNM-like modes $\tilde{\bm{E}}_{n}^{\mathrm{QNM}}(\bm{r})$, which are good approximations of the true QNMs of the system and do not depend on the parameters of the PMLs, and so-called numerical or PML modes $\tilde{\bm{E}}_{n}^{\mathrm{PML}}(\bm{r})$ that depend sensitively on the PML parameters and which originate from the continuous non-resonant contribution in Eq. (5). Hence, by introducing finite-thickness PMLs, this background is effectively mapped to a discrete set of modes olyslager2004discretization. Since the operator in Eq. (II) is not hermitian, the completeness of these eigenmodes is not always guaranteed and we must further assume the absence of exceptional points where two eigenvectors coalesce. In this case, the QNMs and PML modes form a complete set, even in the presence of non-trivial backgrounds, and the PML-regularized scattered field $\tilde{\bm{E}}_{\mathrm{sca}}$ can be expanded in this discrete set of modes,
@@
where, henceforth, we shall not distinguish between the QNM-like and PML modes. Mathematically, they arise as a common set of solutions to the eigenvalue problem in Eq. (II), and should be treated on an equal footing for the purpose of further analysis.
In the context of non-dispersive magnetic permeabilities $\mu=\mu(\bm{r})$, Yan et al. yan2018rigorous have developed an approach for rigorous QNM analysis of lossy resonators, relying on an auxiliary-field eigenvalue technique with PML regularization, within the framework of a finite-element analysis that enables efficient numerical calculations. Their scheme effectively linearizes the non-linear eigenvalue problem embodied by Eq. (II) for dispersive materials, described using an $N$-pole, Drude-Lorentz model:
$\varepsilon(\bm{r},\omega)=\varepsilon_{\infty}(\bm{r})-\varepsilon_{\infty}(\bm{r})\sum_{l=1}^{N}\frac{\omega_{\text{p},l}^{2}(\bm{r})}{\omega^{2}-\omega_{0,l}^{2}(\bm{r})+\mathrm{i}\gamma_{l}(\bm{r})\omega},$ (13)
[…]
@@
$=1\,$, (14)
in which the volume integration is performed over the entire computational domain $\Omega$, including the PML regions $\Omega_{\mathrm{PML}}$, yet the normalization of the QNMs does not depend on the PML parameters sauvan2022norm. At this point, it is worth noting that different normalization procedures have been proposed in the literature, whose validity and equivalence have been the subject of intense debate kristensen2015normalization; muljarov2017comment; kristensen2017reply; lalanne2020mode; for a mathematical and numerical comparison, we refer the interested reader to Ref. sauvan2022norm. For the excitation coefficients in Eq. (II), an analytical form can be derived yan2018rigorous:
@@
$-\frac{\tilde{\omega}_{n}}{\tilde{\omega}_{n}-\omega}\mu_{0}$ $\langle\tilde{\bm{H}}^{*}_{n}|\Delta\mu(\bm{r})|\bm{H}_{\mathrm{b}}\rangle_{L^{2}(\Omega_{\mathrm{res}})}\,.$ (15)
A modal expansion as in Eq. (II) is also possible in the time domain, with coefficients $\xi_{n}(t)$ related to $\alpha_{n}(\omega)$ via a Fourier transform and satisfying the equation
@@
$\hskip-14.22636pt-\varepsilon_{0}\langle\tilde{\bm{E}}_{n}^{*}|\varepsilon_{\mathrm{b}}-\varepsilon_{\infty}|\frac{\mathrm{d}}{\mathrm{d}t}\bm{E}_{\mathrm{b}}(t)\rangle-\mathrm{i}\tilde{\omega}_{n}\mu_{0}\langle\tilde{\bm{H}}_{n}^{*}|\Delta\mu|\bm{H}_{\mathrm{b}}(t)\rangle$
@@
The system Green’s function can also be expanded in the modes according to
with $A_{n}=\varepsilon_{0}c^{2}/\omega(\tilde{\omega}_{n}-\omega)$ or $A_{n}=\varepsilon_{0}c^{2}/\tilde{\omega}_{n}(\tilde{\omega}_{n}-\omega)$, depending on the pole structure of the Green’s function at $\omega=0$ kristensen2020modeling. It is worth noting that although this auxiliary-field eigenvalue approach was originally introduced for systems with non-dispersive magnetic permeabilities $\mu(\bm{r})$, its generalization for media with strongly frequency-dependent permeabilities (as relevant for magnonic cQED) may be feasible via the introduction of suitable auxiliary fields for the magnetic response. In the following, we assume that such a complete set of suitably regularized and normalized modes can be ascertained, and explore their quantization for arbitrary, linear, 3D, magnetodielectric systems.